OpenStudy (anonymous):

How many pounds of an $83 per pound substance must be mixed with an $11 per pound substance to get 240 lbs of an alloy which will cost $28 per pound? Round to the nearest pound.

OpenStudy (anonymous):

83lb + 11 ( 240 - lb) = 28 * 240 ->72lb = 4080 => lb = 57 lbs